T= Tris




A= Glacial acetic acid




E= EDTA




B= Boric Acid




10x TAE Recipe




For 1L of 10x solution,




• 48.5 g tris




• 11.4 mL glacial acetic acid




• 20 mL 0.5M EDTA (pH 8.0)




TBE Buffer 10x Stock Recipe




• 108 g tris base




• 55 g boric acid




• 900 ml double-distilled H2O




• 40 ml 0.5 M EDTA solution (pH 8.0)




Adjust volume to 1 L find molarity

1
Expert's answer
2022-04-26T03:46:02-0400

"n(T)=48.5\/121=0.4 mol; c(T)= 0.4\/1=0.4 mol\/l"

"m(glacial acetic acid) = v*density=11.4*1.0492=11.96 g, c(glacial acetic acid)= 11.96\/(1*60)=0.2 mol\/l"

"c(T)=108\/(121*1)=0.89 mol\/l\nc(boric acid) = 55\/(62*1) = 0.89 mol\/l"